### v1

@@ 6 Performance Evaluation
On-chain per-operation overhead. In this experiment, we evaluate the overhead of the deposit and the syncing process and compare that to the baseline Uniswap on-chain operations. We set the daily volume $V_{D}$ to be $500$K transactions (10x Uniswap). We use a Gas Profiler [22] to measure the gas cost of the different components of the $\mathsf{Sync}$ transaction. We find that storing the state of the liquidity positions is the most expensive, as each consists of 192 bytes (or 6 words), incurring 22,100 Gas units per word. Each payout transaction incurs a constant fee of 15,771. The threshold signature-based quorum certificate incurs a fixed fee that corresponds to the gas cost of the BN256 operations, and a fee that is proportional to the length of the summary data structure. We present an itemized gas cost analysis for the sync and deposit transactions in Table 6.
Table 6: Average latency and itemized gas cost comparison.
@@
—c—l—c—c—c—c—c—c— Transaction Sync   Deposit
[…]
(each) (per 32 byte)  Hash To Point   Verify Pairing
Keccak256 ecMUL
Average Gas 15,771 22,100 $30+6\times\lceil\frac{|\mathsf{summary}|}{256}\rceil$ 6,000 113,000   52,696
Latency(s) 15.28   54.60
[…]
Average gas 160,601.45   435,609.86   158,473.43   163,743.04
Latency (s) 31.34   42.24   12.72   13.45
Table 7: Operation storage overhead comparison.
$\mathsf{ammBoost}$ $\mathsf{Sync}$ component Swap entry Position entry $\mathsf{vk}_{c}$ Signature
@@
Overall, the gas cost of the $\mathsf{Sync}$ call is affected by the number of positions processed in an epoch; this cost does not scale with the number of processed transactions, but rather with the number of users. On the other hand, in baseline Uniswap the gas cost is proportional to the total generated traffic, where the numbers in Table 6 are per one transaction from each type. For the average latency, a $\mathsf{Sync}$ transaction does not depend on any other mainchain transactions, so it is confirmed within one block on average. However, since a deposit depends on two ERC20 approvals, and performs 2 transfers, it takes around 4 blocks. The same behavior is observed in our Uniswap baseline, as a swap requires 1 approval and a mint requires 2 approvals.
@@
We also report the per-operation storage cost. In particular, we report the cost breakdown for the $\mathsf{Sync}$ call the mainchain and the summary-block size for $\mathsf{ammBoost}$, and the transaction sizes for baseline Uniswap on the mainchain. For the $\mathsf{Sync}$ call, the sizes of swap and position entries vary greatly between the summarized changes in a summary-block and the $\mathsf{Sync}$ inputs submitted to the mainchain. This is due to the difference in encoding and binary packing between the sidechain and the mainchain. On the mainchain, Ethereum’s application binary interface (ABI) packing keeps track of the data and all the information needed to reinterpret it back, while on the sidechain we resort to using simple binary packing. We also have an extra 6 words (192 bytes) storage overhead on the mainchain needed for the BLS signature and its public key (namely, $\mathsf{vk}_{c}$) for authenticating the $\mathsf{Sync}$ call. We present our findings in Table 7.
@@
For Uniswap, we notice that the transactions on Sepolia (Table 7) are smaller than the ones we observe on Ethereum (Appendix 0.C). This is because these chains use different Uniswap transaction routers. The calls to the universal router used on Ethereum end up requiring more arguments, resulting in longer transactions. Uniswap on Sepolia deploys a simpler transaction router. Of note is that the simple router contract (Uniswap V3 router) and the more complex of the two (the Universal router) are both available on Ethereum, but the Universal router is not deployed to Sepolia.
@@
Overall comparison to Uniswap V3. We report the total gas cost and the mainchain state growth of the baseline Uniswap and $\mathsf{ammBoost}$. We set the daily volume $V_{D}$ to be $500$K transactions (10x Uniswap) with the default traffic distribution. We measure the overall mainchain gas cost of relevant operations, and the state growth of the mainchain.
Figure 6: Gas cost and chain growth for Uniswap (baseline) and $\mathsf{ammBoost}$.
@@
As shown in Figure 6, even if the sync transactions end up being heavy on gas as the number of positions increases, we achieve a 94.53% gas reduction when compared to Uniswap Sepolia. The high gas cost of the $\mathsf{Sync}$ transaction is offset by it being uncommon (one occurrence per epoch). On the other hand, the gas cost of swaps, mints, burns, and collects in Uniswap are high since all are processed on the mainchain (while in $\mathsf{ammBoost}$ these are processed on the sidechain). A similar trend is observed for the mainchain state growth, where $\mathsf{ammBoost}$ provides 80.25% decrease in growth compared to Uniswap on Sepolia, and 92.80% decrease when compared to Uniswap on production Ethereum. [^6] The growth for Uniswap on production Ethereum is calculated by multiplying the count of each transaction type in our experiment by its size as reported in Appendix 0.C.
@@ 7 Conclusion
We presented $\mathsf{ammBoost}$, a secure state growth controller and throughput booster for AMMs. It combines a new sidechain architecture, called chainBoost, with a functionality split of the AMM. The AMM is divided into two parts: a base smart contract called $\mathsf{TokenBank}$ residing on the mainchain, which manages token pools, users’ deposits and payouts, and any transaction type that must be handled by the mainchain. And a sidechain part that does most of the heavy lifting by processing swaps, mints, burns, and collects. $\mathsf{ammBoost}$ introduces several techniques to address challenges resulting from the unidirectional dependency between the mainchain and the sidechain. We analyze the security of our system and conduct thorough performance evaluation experiments. The results show the great potential of $\mathsf{ammBoost}$ in reducing the on-chain storage footprint of AMMs and boosting their scalability.
@@ Appendix 0.A Security Analysis
In order to prove Theorem 4.1, we prove two lemmas showing that $\mathsf{ammBoost}$ preserves safety and liveness of the underlying AMM (the proofs of these lemmas are inspired by the proofs found in [54]).
@@
Since $\mathsf{ammBoost}$ implements meta-block pruning and sidechain state syncing, we identify the following threats that may impact safety in our system:
  • Invalid processing of AMM transactions: the sidechain committee does not follow the AMM logic in processing transactions, or accept transactions from users who do not own proper deposits on the mainchain, or process these transactions based off an invalid initial state of the token pool.
  • Out-of-sync AMM state on the mainchain: A committee leader does not issue a $\mathsf{Sync}$ function call at the end of the epoch, causing the state of the AMM on the mainchain and the state maintained on the sidechain to be out of sync.
@@
Invalid processing of AMM transactions. The sidechain in $\mathsf{ammBoost}$ uses the same logic adopted by the underlying AMM to process all transactions. This is based off the latest state of the AMM (with respect to user deposits, liquidity positions, and pool balances from $\mathsf{TokenBank}$ on the mainchain). The use of a secure PBFT protocol, with a committee size that guarantees having at maximum $f$ malicious parties (among a committee of size $23f+$ 2) with overwhelming probability as mentioned earlier, guarantees that only valid records that conforms with these rules will be accepted in meta-blocks. Also, this guarantees that summary-blocks are also valid based on the summary rules in $\mathsf{ammBoost}$ and these meta-blocks. Meta-blocks do not get pruned until their $\mathsf{Sync}$ call transaction is confirmed on the mainchain, so anyone can verify their validity as well as the validity of the summary-block.
Out-of-sync AMM state on the mainchain. This is mitigated by the mass-syncing process. An unresponsive leader who does not initiate an agreement on the $\mathsf{Sync}$ function call, or does not submit the result of the agreement to $\mathsf{TokenBank}$, is easily detected by the new committee as no function call is recorded on the mainchain. The new committee then syncs $\mathsf{TokenBank}$ based on the summaries it produces in its epoch and the ones in the previous (one or multiple) epoch. Same for any rollbacks that may happen on the mainchain, mass-syncing will include all summaries that has been lost due to the rollback.
Invalid syncing. If a leader issues invalid syncing information, the committee (which has honest majority) will not endorse these inputs. So simply this invalid syncing information will be ignored, and mass-syncing (discussed above) will handle the syncing within the next epoch. Having an illegitimate committee pretend to be the rightful one to submit (invalid) sync is addressed in $\mathsf{ammBoost}$ using the syncing authentication mechanism discussed in Section 4.2. The committee of epoch $e$ will not accept the generated committee public verification key $\mathsf{vk}_{c}$ unless there are valid proofs of election showing that the newly claimed committee is the rightful one (so this relies on the security of the election mechanism to be publicly verifiable). Furthermore, an illegitimate committee instead may try to forge a signature over the syncing information under a valid $\mathsf{vk}_{c}$, which succeeds with negligible probability by the security of the used threshold digital signature scheme.
@@
Accordingly, $\mathsf{ammBoost}$ satisfies the safety of the AMM state and will not result in any correctness or security threats.∎
@@
  • Violating the public verifiability of the sidechain: this covers all threats related to the syncing and pruning of meta-blocks that may impact the public verifiability of the sidechain (which in turn impacts the public verifiability of the AMM state).
Proving that $\mathsf{ammBoost}$ mitigates these threats is the same as in [54]. For completeness, we review that proof arguments here. DoS is addressed by having a rotating committee election (a new committee is elected for each new epoch) such that this committee have an honest majority. A leader that targets particular users, and so omit their transactions from all proposed meta-blocks, will operate for one epoch and then a new committee with a new leader will be elected for the next epoch. Thus, maintaining a situation where all future leaders are malicious and perform the same DoS is unfeasible. Sidechain liveness is satisfied due to the use of a secure PBFT consensus and having a large enough committee that satisfies honest majority (so they will be active during the agreement). Also, leader-change allows changing a leader who deliberately attempt to stall the network by not proposing new blocks. Public verifiability is guaranteed by the security of the PBFT consensus; meta-blocks are not pruned until their corresponding $\mathsf{Sync}$ call is confirmed, and summary-blocks are permanent. Also, by having the AMM base smart contract $\mathsf{TokenBank}$ on the mainchain synced correctly based on the sidechain summaries (as discussed in the proof of Lemma 1), summaries are not lost. All of these allow anyone to verify the validity of the evolving state of the AMM. Accordingly, $\mathsf{ammBoost}$ preserves the liveness of the AMM.∎

### v3

@@ 6.2 Comparison with the Baseline
On-chain (itemized) per-operation overhead. We evaluate the overhead of the deposit and the syncing processes in $\mathsf{ammBoost}$, and compare that to the baseline Uniswap on-chain operations. We set the daily volume $V_{D}$ to be $500$K transactions (10x Uniswap). We use a Gas Profiler [23] to measure the gas cost of the different components of the $\mathsf{Sync}$ transaction. As shown in Table 6.2, we find that storing the state of the liquidity positions is the most expensive, as each consists of 192 bytes (or 6 words), incurring 22,100 gas units per word. The same gas cost of 22,100 gas unit per word is incurred when storing the liquidity pool balance. Each payout transaction incurs a constant fee of 15,771 units. A deposit of two tokens incurs a total cost of 105,392 units. The threshold signature-based quorum certificate incurs a fixed fee that corresponds to the gas cost of the BN256 operations needed for verification, and a fee that is proportional to the length of the summary data structure.
Table 1: Mainchain latency and itemized gas cost for $\mathsf{ammBoost}$ operations ($|sum|$ is the size of the summaries).
@@
—l—c—c—c—c—c—c— Operation Sync   De-
Module Payout Storage   Authentication   posit
(each) (per  Hash To Point   Verify   (2 tokens)
32 byte) Keccak256 ecMUL Pairing
Avg. gas 15,771 22,100 $30+6\times\lceil\frac{|sum|}{256}\rceil$ 6,000 113,000   105,392
MC. lat. (s) 15.28   54.60
Table 2: Mainchain latency and gas cost for Uniswap.
Operation Swap Mint Burn Collect
Avg. gas 160,601.45 435,609.86 158,473.43 163,743.04
MC. lat. (s) 31.34 42.24 12.72 13.45
Table 3: Operation storage overhead.
$\mathsf{ammBoost}$ Payout Position $\mathsf{vk}_{c}$ Signature
$\mathsf{Sync}$ component entry entry
@@
Overall, the gas cost of the $\mathsf{Sync}$ call is affected by the number of positions processed in an epoch, the number of deposits made for an epoch, and the updated pool liquidity balance; this cost does not scale with the number of processed transactions, but rather with the number of clients and liquidity providers. On the other hand, in baseline Uniswap the gas cost is proportional to the total generated traffic, where the numbers in Table 2 are per one transaction from each type. For the average latency, a $\mathsf{Sync}$ transaction does not depend on any other mainchain transactions, so it is confirmed within one block on average. However, since a two-token deposit depends on two ERC20 approvals, and performs 2 transfers, it takes around 4 blocks in our experiments. The same behavior is observed in our Uniswap baseline, as a swap requires 1 approval from the user and a mint requires 2 approvals. Since a two-token deposit depends on two ERC20 approvals, and performs two transfers, it will take at least four rounds if the operations are done sequentially. For our Uniswap baseline, as a swap requires one approval, it takes a minimum of two rounds to be processed; a mint takes three rounds at least, at it requires two approvals.
@@
We also report the per-operation storage cost. In particular, we report the cost breakdown for the $\mathsf{Sync}$ call on the mainchain and the summary-block size for $\mathsf{ammBoost}$, and the transaction sizes for baseline Uniswap on the mainchain. For the $\mathsf{Sync}$ call, the sizes of payout and position entries vary greatly between the summarized changes in a summary-block and the $\mathsf{Sync}$ inputs submitted to the mainchain. This is due to the difference in encoding and binary packing between the side and main chains. On the mainchain, Ethereum’s application binary interface (ABI) packing keeps track of the data and all the information needed to reinterpret it back, while on the sidechain we use simple binary packing. We also have an extra 6 words (192 bytes) storage overhead on the mainchain needed for the BLS signature and its public key (namely, $\mathsf{vk}_{c}$) for authenticating the $\mathsf{Sync}$ call. We present our findings in Table 3.
@@
As shown, for Uniswap, we notice that the transactions on Sepolia are smaller than the ones we observe on Ethereum (Appendix 0.C). This is because these chains use different Uniswap transaction routers. The calls to the universal router used on Ethereum end up requiring more arguments, resulting in longer transactions. Uniswap on Sepolia deploys a simpler transaction router. Of note is that the simple router contract (Uniswap V3 router) and the more complex of the two (the Universal router) are both available on Ethereum, but the Universal router is not deployed to Sepolia.
Figure 6: Gas cost and chain growth comparison.
In general, and as will discussed next, Uniswap incurs a larger storage cost as its transactions are quite large and all are logged on the mainchain. While for $\mathsf{ammBoost}$, only (the less frequent) $\mathsf{Sync}$ call transaction is logged on the mainchain.
@@
Overall comparison. We report the total gas cost and the mainchain state growth of the baseline Uniswap and $\mathsf{ammBoost}$. We set the daily volume $V_{D}$ to be $500$K transactions (10x Uniswap) with the default traffic distribution. We measure the overall mainchain gas cost of relevant operations, and the state growth of the mainchain.
@@
As shown in Figure 6, even if the sync transactions end up being heavy on gas as the number of positions and payouts increases, we achieve a 96.05% gas reduction when compared to Uniswap Sepolia. The high gas cost of the $\mathsf{Sync}$ transaction is offset by it being uncommon (one occurrence per epoch). On the other hand, the gas cost of swaps, mints, burns, and collects in Uniswap are high since all are processed on the mainchain (while in $\mathsf{ammBoost}$ these are processed on the sidechain). A similar trend is observed for the mainchain state growth, where $\mathsf{ammBoost}$ provides 93.42% decrease in growth compared to Uniswap on Sepolia, and 97.60% decrease when compared to Uniswap on production Ethereum. [^15] The growth for Uniswap on production Ethereum is calculated by multiplying the count of each transaction type in our experiment by its size as reported in Appendix 0.C
6.3 Impact of Parameter Configuration
We study the impact of parameter configuration on $\mathsf{ammBoost}$’s performance, including traffic amount and distribution, block size, sidechain round duration, and number of rounds per epoch.
Scalability. In this experiment, we test the scalability of $\mathsf{ammBoost}$ (for a single pool) to understand its behavior under heavy traffic. We follow the same traffic distribution as in Uniswap and vary the daily volume $V_{D}\in\{50K,500K,5M,25M\}$. We record the impact on throughput and transaction/payout latency as shown in Table 4.
Table 4: Scalability of $\mathsf{ammBoost}$.
Daily volume 50K 500K 5M 25M
Throughput (tx/s) 0.42 3.41 33.04 138.06
Avg. sc latency (s) 7.13 7.13 7.13 231.52
Avg. payout latency (s) 120.71 120.71 120.71 346.49
Throughput-wise, we record a low throughput of 0.42 tx/s to 33.04 tx/s for a daily volume of 50K to 5M transactions (roughly 1x-100x Uniswap’s daily volume). This is mainly due to the mainchain blocks not being full as this workload is way below the capacity that $\mathsf{ammBoost}$ can handle. While for traffic that is 500x Uniswap’s daily volume, $\mathsf{ammBoost}$ achieves a throughput of 138.06 tx/s.
Latency-wise, we achieve a quasi-instant when the daily volume is between 50K and 5M (transactions that arrive at the beginning of the round get processed within the same round, while the residual amount of latency is due to transactions generated at the end of the epoch and processed in the next epoch). This leads to payouts being processed within one epoch (including the time needed to confirm the $\mathsf{Sync}$ transaction on the mainchain). Transaction congestion happens when the daily volume is 25M, resulting in higher average transaction and payout latency as the table shows.
Impact of block size. We test the impact of the sidechain block size with the goal of finding an optimal block size for our system. Thus, we compare different deployments of $\mathsf{ammBoost}$ with different block sizes against Uniswap’s Sepolia deployment. We run the protocol with the following block sizes {$0.5,1,1.5,2$} MB, and we increase the daily volume to 50M transactions. We measure the impact on throughput and transaction/payout latency with the goal of identifying the block size that maximizes throughput while minimizing latency. Our results can be found in Table 5.
Table 5: Impact of different sidechain block sizes.
BlockSize (MB) 0.5 1 1.5 2
Throughput (tx/s) 68.97 138.61 207.52 276.43
Avg. sc. latency (s) 4357.00 1603.01 687.98 230.48
Avg. payout latency (s) 4472.63 1719.10 804.05 345.44
As expected, increasing the block size improves both throughput and latency, as more transactions can be packed in a block which reduces queue congestion. However, larger block sizes mean a larger propagation delay that could be problematic for short sidechain round duration. Thus, system designers should be careful when choosing an optimal sidechain block size, balancing between the block size that can handle the daily volume of transactions while capturing the intricacies of large network transfers.
Impact of sidechain round duration. Another important factor to study is the impact of the sidechain round duration. An ideal round duration should allow for consensus to conclude while maximizing throughput and minimizing latency. As PBFT agreement takes on average around 6 sec to conclude in our implementation, we test the following round duration values: 7, 11, 16, and 21 seconds, and report the performance metrics as before (Table 6).
Table 6: Impact of different sidechain round durations.
Sc round duration (s) 7 11 16 21
Throughput (tx/s) 138.06 92.18 61.75 46.31
Avg. sc latency (s) 231.52 921.64 1950.92 2975.90
Payout latency (s) 346.49 1087.95 2193.85 3295.11
Throughput-wise, we observe that as the block time increases, throughput decreases and latency increases. This is due to processing the same amount of transactions while increasing the time needed to produce a block. To choose an optimal block time, system designers need to take into consideration the time required for the sidechain consensus and network propagation delays, while aiming to generate new blocks as fast as possible.
Impact of the number of sidechain rounds per epoch. We test the impact of the number of sidechain rounds within an epoch. The goal is to find an epoch length that maximizes throughput, and minimizes transaction/payout latency, based on the optimal sidechain round duration from the experiment above. Thus, we pick our epoch to have $\{5,10,20,30,60,96\}$ sidechain rounds, each of which lasts 7 sec, and report the performance metrics (Table 7).
Having short epochs negatively affects throughput and the sidechain latency. As a matter of fact, frequent summary-blocks harm performance since this leads to a larger number of $\mathsf{Sync}$ calls that are costly. At the same time, fewer transactions are processed within the epoch, thus affecting both latency and throughput. Longer epoch duration reduces the sidechain latency and increases throughput. However, this affects the payout latency adversely since $\mathsf{Sync}$ calls now are much fewer and spaced out (so users have to wait longer, as the epoch itself is longer, to obtain their actual token payouts). Also, this means that these users have to put larger deposits to cover their (long) epoch activities, which could be undesirable. Based on our results, we achieve the best payout latency when the epoch lasts for 20 sidechain rounds, which is equivalent to 140 sec.
Impact of traffic distribution. In this experiment, we evaluate different traffic distributions as follows (all the numbers are percentages): $(s,m,b,c)\in\{(60,20,10,10)$, $(60,10,20,10),$ $(60,10,10,20),$ $(80,10,5,5),$ $(80,5,10,5),$ (80,5,5,10)}(80,5,5,10)\}( 80 , 5 , 5 , 10 ) }, where $(s,m,b,c)$ stand for swaps, mints, burns, and collects, respectively. As noted, in these configurations we keep the swap operations dominant to align with the baseline AMM traffic distribution observed in practice (see Appendix 0.C). Our results can be found in Table 8.
Table 7: Impact of number of sidechain rounds per epoch.
Epoch len 5 10 20 30 60 96
(sc rounds)
Throughput 114.27 128.53 135.90 138.06 140.66 141.53
(tx/s)
SC latency (s) 517.94 333.54 255.57 231.52 208.96 199.55
Payout 545.12 337.86 334.81 346.49 434.94 546.04
latency (s)
Table 8: Impact of traffic distribution.
Swap % 60 80
Mint % 20 10 10 10 5 5
Burn % 10 20 10 5 10 5
Collect % 10 10 20 5 5 10
Throughput 145.16 143.76 140.91 143.76 140.23 140.14
(tx/s)
SC latency (s) 162.26 175.35 177.39 202.48 215.06 210.35
Payout 277.99 291.05 293.03 317.23 329.81 324.43
latency (s)
Max sc 31831 31831 31831 31831 31831 31831
growth (B)
@@ 7 Conclusion
We presented $\mathsf{ammBoost}$, a secure state growth controller and throughput booster for AMMs. It combines a dependent-sidechain architecture with a functionality split of the AMM. The AMM is divided into two parts: a base smart contract called $\mathsf{TokenBank}$ residing on the mainchain, which manages token pools, users’ deposits and payouts, and any transaction type that must be handled by the mainchain. And a sidechain part that process most of the workload. $\mathsf{ammBoost}$ introduces several techniques to address challenges resulting from the unidirectional dependency between the mainchain and the sidechain. We analyze the security of our system and conduct thorough performance evaluation experiments. The results show the great potential of $\mathsf{ammBoost}$ in reducing the on-chain storage footprint of AMMs and boosting their scalability.
Acknowledgments
The work of M.E.N. is supported by NSF under Grant No. CNS-2226932, and the work of G.A. is supported by the Latest in DeFi Research (TLDR) fellowship funded by Uniswap Foundation.
@@ Appendix 0.A Security Analysis
In order to prove Theorem 4.1, we prove two lemmas showing that $\mathsf{ammBoost}$ preserves safety and liveness of the underlying AMM (these proofs are inspired by those found in [57]).
@@
Since $\mathsf{ammBoost}$ implements meta-block pruning and mainchain (i.e., $\mathsf{TokenBank}$) state syncing, we identify the following threats that may impact safety in our system:
  • Invalid processing of AMM transactions: the sidechain committee does not follow the AMM logic in processing transactions, or accept transactions from users who do not own enough deposits on the mainchain, or process these transactions based off an invalid initial state of the token pool.
  • Out-of-sync AMM state on the mainchain: A committee leader does not issue a $\mathsf{Sync}$ function call at the end of the epoch, causing the state of the AMM on the mainchain and the state maintained on the sidechain to be out of sync. This also could happen due to rollbacks on the mainchain causing recent $\mathsf{Sync}$ calls to be lost.
@@
Invalid processing of AMM transactions. The sidechain in $\mathsf{ammBoost}$ uses the same logic adopted by the underlying AMM to process all transactions. This is based off the latest state of the AMM (with respect to user deposits, liquidity positions, and pool balances from $\mathsf{TokenBank}$ on the mainchain). The use of a secure PPFT protocol, with a committee size that guarantees having at maximum $f$ malicious parties (among a committee of size $23f+$ 2) with overwhelming probability as mentioned earlier, guarantees that only valid records that conform with these rules will be accepted in meta-blocks. Also, this guarantees that summary-blocks are also valid based on the summary rules in $\mathsf{ammBoost}$ and these meta-blocks. Meta-blocks do not get pruned until their $\mathsf{Sync}$ call transaction is confirmed on the mainchain, so anyone can verify the validity of these blocks and the validity of the summary-block (as well as the $\mathsf{Sync}$ call).
Out-of-sync AMM state on the mainchain. This is mitigated by the mass-syncing process. An unresponsive leader who does not initiate an agreement on the $\mathsf{Sync}$ function call, or does not submit the result of the agreement to $\mathsf{TokenBank}$, is easily detected by the new committee as no function call has been issued and recorded on the mainchain. The new committee then syncs $\mathsf{TokenBank}$ based on the summaries it produces in its epoch and the ones in the previous (one or multiple) epoch. Same for any rollbacks that may happen on the mainchain, mass-syncing will include all summaries that has been lost due to the rollback.
Invalid syncing. If a leader issues invalid syncing information, the committee (which has honest majority) will not endorse these inputs. So simply this invalid syncing information will be ignored, and mass-syncing (discussed above) will handle the syncing within the next epoch. Having an illegitimate committee pretend to be the rightful one to submit (invalid) sync is addressed in $\mathsf{ammBoost}$ using the syncing authentication mechanism discussed in Section 4.2. The committee of epoch $e$ will not accept the generated committee public verification key $\mathsf{vk}_{c}$ unless there are valid proofs of election showing that the newly claimed committee is the rightful one (so this relies on the security of the election mechanism to be publicly verifiable). Furthermore, an illegitimate committee (or an attacker) instead may try to forge a signature over the syncing information under a valid $\mathsf{vk}_{c}$, which succeeds with negligible probability by the security of the used threshold digital signature scheme.
@@
Accordingly, $\mathsf{ammBoost}$ satisfies the safety of the AMM and will not result in any correctness or security threats.
@@
  • Violating the public verifiability of the sidechain: this covers all threats related to the syncing and pruning of meta-blocks that may impact the public verifiability of the sidechain (which in turn impacts the public verifiability of the AMM).
Proving that $\mathsf{ammBoost}$ mitigates these threats is the same as in [57]. For completeness, we review that proof arguments here. DoS is addressed by having a rotating committee election (a new committee is elected for each new epoch) such that this committee has an honest majority. A leader that targets particular users, and so omit their transactions from all proposed meta-blocks, will operate for one epoch and then a new committee with a new leader will be elected for the next epoch. Thus, maintaining a situation where all future leaders are malicious and perform the same DoS is unfeasible. Sidechain liveness is satisfied due to the use of a secure PBFT consensus and having a large enough committee that satisfies honest majority (so they will be active during the agreement). Also, leader-change allows changing a leader who deliberately attempt to stall the network by not proposing new blocks, and mass-syncing will address the case of a malicious leader who does not initiate agreement on a summary-block or $\mathsf{Sync}$ function call. Public verifiability is guaranteed by the security of the PBFT consensus; meta-blocks are not pruned until their corresponding $\mathsf{Sync}$ call is confirmed, and summary-blocks are permanent. Also, by having the AMM base smart contract $\mathsf{TokenBank}$ on the mainchain synced correctly based on the sidechain summaries (as discussed in the proof of Lemma 1), summaries are not lost. All of these allow anyone to verify the validity of the evolving state of the AMM. Accordingly, $\mathsf{ammBoost}$ preserves the liveness of the AMM.